COLLISION-INDUCED ABSORPTION IN PURE O_{2}, CO_{2}, AND N_{2}

\underline{Y. I. BARANOV}, W. J. LAFFERTY AND G. T. FRASER, Optical Technology Division, NIST, Gaithersburg, MD 20899.

Collision-induced absorption (CIA) of the fundamental band of pure oxygen has been obtained over the temperature range -80^{o} to +85^{o}C. The integrated band intensity is at a minimum at 0^{o}C and increases on either side of this temperature. Band intensities of the CIA in the fundamental band of N_{2} as well as the CO_{2} Fermi dyad are currently under investigation. Preliminary results indicate that the temperature dependence of the band intensity of the nitrogen CIA parallels that of O_{2}. In contrast, the integrated intensities of the 10^{0}0/02^{0}0 Fermi dyad in CO_{2} increase steadily as the temperature is decreased. The integrated intensities double as the temperature goes from +72^{o} to -80^{o} C. This indicates that at low temperatures the bound dimer is the major species in the gas. The band structure obtained at elevated temperatures will be presented.